SkyChefs, Inc., prepares in-flight meals for a number of major airlines. One of the company’s products is grilled salmon in dill sauce with baby new potatoes and spring vegetables. During the most recent week, the company prepared 4,000 of these meals using 750 direct labor-hours. The company paid its direct labor workers a total of $6,000 for this work, or $8.00 per hour. According to the standard cost card for this meal, it should require 0.20 direct labor-hours at a cost of $7.00 per hour.

Required:

1. What is the standard labor-hours allowed (SH) to prepare 4,000 meals?

2. What is the standard labor cost allowed (SH × SR) to prepare 4,000 meals?

3. What is the labor spending variance?

4. What is the labor rate variance and the labor efficiency variance?
Answer #1
Answer
1) Standard labour hour allowed = 4000*0.20 = 800 hour
2) Standard cost allowed = 800*7 = 5000
3) Labour spending variance = 5000-6000 = 1000 A
4) Labour rate variance = (7*750-6000) = 750 U
Labour efficiency variance = (800-750)*7 = 350 F
Please Like ( Positive Rating )